Extra: Let X, Y, Z be results of three independent tosses of a fair die. (a) Find the covariance of the random variables W=2X-3Y + Z (b) Find the correlation coefficient of W and V. and V=X-2Y-Z
Answer #1

(a)

For a fair coin, p = 1/2

Var(X) = Var(Y) = Var(Z) = p(1-p) = (1/2)(1 - 1/2) = 1/4

Cov(W, V) = Cov(2X - 3Y + Z, X - 2Y - Z)

= Cov(2X, X - 2Y - Z) + Cov(- 3Y, X - 2Y - Z) + Cov(Z, X - 2Y - Z)

= Cov(2X, X) + Cov(2X, - 2Y) + Cov(2X, - Z)

+ Cov(- 3Y, X) + Cov(- 3Y, - 2Y) + Cov(- 3Y, - Z)

+ Cov(Z, X) + Cov(Z, - 2Y) + Cov(Z, - Z)

= 2Cov(X, X) - 2*2 Cov(X, Y) + 2 * (-1) Cov(X, Z)

+ (-3)*Cov(Y, X) + (-3) * (-2)Cov(Y, Y) + (-3)*(-1)Cov(Y, Z)

+ Cov(Z, X) + (-2)Cov(Z, Y) + (-1)Cov(Z, Z)

= 2 Var(X) + 6 Var(Y) - Var(Z)

= 2 * (1/4) + 6 * (1/4) - (1/4)

= 7/4

(b)

Var(W) = Var(2X - 3Y + Z) = 22Var(X) + (-3)2Var(Y) + Var(Z)

= 4 * (1/4) + 9 * (1/4) + (1/4) = 14/4 = 7/2

Var(V) = Var(X - 2Y - Z) = Var(X) + (-2)2Var(Y) + (-1)2Var(Z)

= (1/4) + 4 * (1/4) + (1/4) = 6/4 = 3/2

Corr(W, V) = Cov(W, V) / Var(X)Var(Y)

= (7/4) / sqrt{(7/2)*(3/2)}

= 0.7638